in elmination how would i solve this:x+2y=-9
4x+2y=6

Answer by nerdybill(7384)

x+2y=-9
4x+2y=6
.
Multiply both sides of top equation by -1:
-x-2y=9
4x+2y=6
.
Now, add the two equation together:
-x-2y=9
4x+2y=6
-------
3x =18
x = 6 (solution for x)
.
Substitute the above back into the original top equation and solve for y:
x+2y=-9
6+2y=-9
2y=-15
y=-15/2 (solution for y)